OpenStudy (anonymous):

The sum of two numbers is 57. One of the numbers is 21. Which of the following equations could be used to find the other number? 57 = n + 21 n = 57 + 21 57 = 2 + 21 Not Here 21 = n – 57

OpenStudy (agreene):

57 = n + 21

Directrix (directrix):

57 - 21 = n

Directrix (directrix):

This will work: 57 = n + 21

OpenStudy (anonymous):

57=n+21